# Agentic Large Language Models, a survey

### v2

@@ 1.5 Taxonomy
| Toolformer (Schick et al., 2023) | Call APIs | Tool calling | Calculator, Search engine |
| ToolBench (Qin et al., 2023) | 16,464 APIs | Tool calling | API framework |
| EasyTool (Yuan et al., 2024b) | Tool documentation | Tool calling | ToolBench |
| ToolAlpaca (Tang et al., 2023) | 400 APIs | Tool calling | Benchmark |
| ToolQA (Zhuang et al., 2023) | APIs | Tool calling | Question answering |
| Gorilla (Patil et al., 2023) | Generate APIs | Tool calling | APIBench |
| AgentHarm (Andriushchenko et al., 2024) | Adversarial Agents | Robust LLMs | Adversarial Benchmark |
@@
| RainbowTeaming (Samvelyan et al., 2024) | MAP Elites | Robust LLMs | Adversarial Benchmark |
@@
| AssistantGPT (Neszlényi et al., 2024) | Websearch, OpenAPI, Voice | Tools, Planner, Memory | Education/Corporate |
| Meeting Assist (Cabrero-Daniel et al., 2024) | LLM | Meetings | Scrum |
@@
| MUCA (Mao et al., 2024) | topic generator | What/When/How | Group Conversations |
| Task Scheduling (Bastola et al., 2023) | LLM | Task Scheduling | Collaborative Group |
| Thinking Assistant (Park and Kulkarni, 2023) | LLM | Human reflection | Human |
| LLaSa (Zhang et al., 2024c) | finetuned LLM, CoT, RAG | E-commerce assistant | ShopBench |
| MMLU (Jin et al., 2024b) | shopping skills | Finetuning | Benchmark |
| Question suggestion (Vedula et al., 2024) | LLM | Product metadata | Shopping |
| ChatShop (Chen et al., 2024a) | finetuned LLM | Information-seeking | Shopping |
| Flight Booking Assistant (Manasa et al., 2024) | finetuned LLM, RAG | Flight Booking | Booking process |
@@
| Medical Note generation (Yuan et al., 2024a) | finetuned LLM | Medical Scribe | Medical note taking |
@@
| Medical Reports (Sudarshan et al., 2024) | Reflexion (Shinn et al., 2024) | 21st Century Cures Act | Health records |
| MedCo (Wei et al., 2024a) | Multiagent Copilot | Medical education | education |
| Benchmark (Qiao et al., 2024) | RAG | Agentic Workflow | Benchmark |
| Wind Hazards (Tabrizian et al., 2024) | LLM | Flight Planning | Flight Operations |
| Flight Dispatch (Wassim et al., 2024) | LLM | Drone as a Service | Flight Operations |
| FinAgent (Zhang et al., 2024b) | Multimodal, RAG | Analysis modules | Stock data |
| FinRobot (Yang et al., 2024a) | finetuned LLM | Document Analysis | Financial Documents |
| FinMem (Yu et al., 2024b) | Multi-agent | Trading Agent Assistant | Market data |
| TradingAgents (Xiao et al., 2024) | Multi-agent | Collaborative dynamics | Simulation |
| AI Scientist (Lu et al., 2024a) | Chain of Thought | Reflexion (Shinn et al., 2024) | Scientific experiment |
| SWE-Agent (Yang et al., 2024b) | Codex | ReAct (Yao et al., 2022) | Agent-Computer Interface |
| MLGym (Nathani et al., 2025) | Chain of Thought | SWE-Agent (Yang et al., 2024b) | Gym (Brockman et al., 2016) |
@@
| Iterated Prisoner’s (Fontana et al., 2024) | LLM | Cooperate/Defect | Social Dilemma |
@@
| Social Games (Akata et al., 2023) | LLM | Cooperate/Defect | Battle of the Sexes, etc |
@@
| GTBench (Duan et al., 2024) | CoT/ToT | Cooperate/Defect | Kuhn poker, liar’s dice, nim |
@@
| GAMA-Bench (Huang et al., 2024a) | LLM | Cooperate/Defect | El Farol, Public Goods, etc |
| Theory of Mind (van Duijn et al., 2023) | LLM | Theory of Mind | Stories |
| NegotiationArena (Bianchi et al., 2024) | LLM | Dialogue | Negotiation |
| Alympics (Mao et al., 2023) | LLM | Multi-agent sandbox | Water-allocation challenge |
| MAgIC (Xu et al., 2024a) | LLM | social interaction | Social Deduction games |
| AucArena (Chen et al., 2023a) | LLM | Bidding/Goal | Auction |
| EgoSocialArena (Hou et al., 2024) | LLM | Social Intelligence | Cognitive, Situational, Behavioral |
| Donor Game (Vallinder and Hughes, 2024) | LLM | Reciprocity | Social skill Game |
| Social Simulacra (Park et al., 2022) | LLM | Society | Simulation of Society, Party |
| Reconcile (Chen et al., 2023b) | LLM | Concensus | Round Table Conference |
@@
| MindStorms (Zhuge et al., 2023) | LLM | Society of Mind (Minsky, 1988) | Multi-agent problem solving |
| AutoGen (Wu et al., 2023) | LLM infrastructure | aegnt-agent conversation | Framework |
| AgentVerse (Chen et al., 2023c) | LLM | Group dynamics | Collaborative problem solving |
| ChatEval (Chan et al., 2023) | LLM | Collaborative problem solving | Text summarization |
| CAMEL (Li et al., 2023a) | LLM infrastructure | Multi-agent interaction | Roleplaying Framework |
| OASIS (Yang et al., 2024e) | lightweight LLM | Social media simulator | Reddit/X |
| WebArena (Zhou et al., 2023a) | Web benchmark | e-commerce, forum, content | Benchmark |
| Balrog (Paglieri et al., 2024) | RL games | interaction | Benchmark |
| BenchAgents (Butt et al., 2024) | Planning | human in the loop | Benchmark |
| AgentBoard (Ma et al., 2024a) | Embodied, Web, Tool | interactions | Benchmark |
| Bias (Fernando et al., 2024) | LLM | healthcare, justice, business | Benchmark |
@@
| Citing (Feng et al., 2023) | Curriculum Learning | Teacher/Student | Instruction Tuning |
| WEBRL (Qi et al., 2024) | Curriculum Learning | Self-evolving | WebArena |
@@
| Expert Iteration (Zhao et al., 2024b) | Curriculum Learning | Reasoning | Hallucination Mitigation |
| EvolutionaryAgent (Li et al., 2024b) | Evolutionary LLM | Norm Aligment | Multi-agent Infrastructure |
@@
| Social Conventions (Ashery et al., 2024) | Naming Game | Norm emergence | Naming game (Steels, 1995) |
| MetaNorms (Horiguchi et al., 2024) | LLM | Norm emergence | Metanorms (Axelrod, 1986) |
| Norm Violations (He et al., 2024) | LLM | Norm violations | 80 household stories |
| CASA (Qiu et al., 2024a) | LLM | Cultural and Social awareness | Benchmark |
| Collaboration (Zhang et al., 2023a) | LLM | 4-traits, cooperation | LLM societies |
@@
| Power hierarchy (Campedelli et al., 2024) | LLM | persuasive/abusive behavior | Stanf Prison Exper (Zimbardo, 1972) |
| Argumentation (Van Der Meer et al., 2024) | Hybrid LLM | LLM supported Argumentation | Benchmark |
@@
| Debate (Baltaji et al., 2024) | LLM-agents | collaboration, debate | Multi-agent discussion |
2 Reasoning
@@
Note that both retrieval augmentation and self reflection can be used to generate new training data. Retrieval augmentation can be used to retrieve relevant information beyond the originally available training dataset. Self reflection uses methods related to planning that imagine plausible futures, that can be useful for training of LLMs.
Originally, the methods for intelligent decision making that we review in this section were developed with the goal of improving the predictive modeling performance of the LLM. For the field of agentic LLMs, the reasoning techniques are used as an important fundament for agents that act with the world, and interact with each other.
2.1 Multi Step Reasoning
@@
We will start by reviewing works that apply reasoning methods to improve decision making, inspired by Chain of Thought’s step-by-step approach (Wei et al., 2022b).
2.1.1 Chain of Thought Step-by-Step
@@
Originally, LLMs performed poorly on math word problems, even on simple grade school math word problems (GSM8K, Cobbe et al. (2021)). LLMs are trained to produce an immediate answer to a prompt, and they typically take shortcuts that may look good, but are semantically wrong.222What is the correct answer to: This is as simple as two minus two is …? The phrase: as simple as two plus two is four may well have a higher frequency in a training corpus than the phrase: as simple as two minus two is zero.
@@
To correctly solve complex reasoning problems, humans are taught to use a step-by-step approach. If a reasoning problem is better solved by following a step-by-step approach, then a sensible approach is to prompt the model to follow suitable intermediate steps, answer those, and work towards the final answer. Wei et al. (2022b) showed in their Chain of Thought paper that with the right prompt the LLM follows such intermediate steps. When the LLM is prompted to first rephrase information from the question as intermediate reasoning steps in its answer, the LLM performed much better than when it was prompted to answer a math problem directly, without reproducing the information from the question in its answer (see their example in Figure 4).
@@
Kojima et al. (2022) find that the addition of a single standard phrase to the prompt (Let’s think step by step) already significantly improves performance. Chain of Thought prompts have been shown to significantly improve performance on benchmarks that included arithmetic, symbolic, and logical reasoning.
@@
Long reasoning chains, however, introduce a challenge, since with more steps hallucination increases. A verification method is needed to prevent error-accumulation. A popular approach is Self Consistency (Wang et al., 2022). Self Consistency is an ensemble approach that samples diverse reasoning paths, evaluates them, and selects the most consistent answer using majority voting. It improves the performance of Chain of Thought typically by 10-20 percentage points when tested on bechmarks. Prompt improvement approaches based on Chain of Thought and Self Consistency are being used in most modern reasoning LLMs, including OpenAI o1, o3 and DeepSeek (Wu et al., 2024; Guo et al., 2025).
2.1.2 Interpreter and Debugger
@@
To solve problems that require mathematical or formal reasoning, it is often advantageous to reformulate the problem into a mathematical or programming language. This translated problem can then be solved by a specialized system, such as a mathematical reasoner, an interpreter, or a planner.
@@
LLMs are not just successful in natural languages, but also in formal (computer) languages. Codex is an LLM that is pretrained on computer programs from GitHub (Chen et al., 2021), which has been successfully deployed. Codex has been used as the basis for the MathPrompter system(Imani et al., 2023). MathPrompter is an ensemble approach that generates algebraic expressions or Python codes, that are then solved using a math solver, or a Python interpreter. Using this approach, MathPrompter achieves state-of-the-art results on the MultiArith dataset (from 78.7% to 92.5%), with GPT-3.
@@
Two other approaches that use a formal language are Program of Thought (PoT) (Chen et al., 2022) and Program Aided Language (PAL) (Gao et al., 2023b). Both approaches generate Python code and use the Python interpreter to evaluate the result.
@@
Debuggers can be used to provide feedback on generated code. This approach is followed in the Self Debugging work (Chen et al., 2023d), that teaches an LLM to debug its generated program code. It follows the same steps of code generation, code execution, and code explanation that a human programmer follows. Several works use Self Debugging to generate code tuned for solving specific problems automatically, without human feedback. Romera-Paredes et al. (2024) introduced FunSearch, an approach that integrates formal methods and LLMs to enhance mathematical reasoning and code generation. It uses a genetic algorithm approach with multiple populations of candidate solutions (programs), which are automatically evaluated (using tools depending on the problem specification). LLaMEA (Large Language Model Evolutionary Algorithm) leverages evolutionary computation methods to generate and optimize evolutionary algorithms (van Stein and Bäck, 2024).
@@
Planners are also combined with LLMs at the language level. Bohnet et al. (2024) provide a benchmark for PDDL (Aeronautiques et al., 1998) based planning problems. They study how LLMs can achieve success in the planning domain (Figure 5).
@@
In the section on retrieval augmentation, we will see further approaches where deep learning and symbolic approaches are successfully combined (Gao et al., 2023b).
2.1.3 Search Tree
@@
Chain of Thought uses a prompt that causes the model to perform a sequence of steps. When there is a single next step, that will be taken. When there are more possibilities, it is unclear how the next step should be selected. A greedy method selects the single step that looks best, follows only that step, and forgets the alternatives (Chain of Thought). Ideally, we should follow the tree of all possible steps. This method is chosen in the Tree of Thoughts approach (Yao et al., 2024). Here, an external control algorithm is created, that calls the model, each time with a different prompt, so that it follows a tree of reasoning steps. When one reasoning path has been traversed, the search backtracks, and tries an alternative. The paper describes both a breadth-first and a depth-first controller.
@@
Together, the trio that consists of a generation prompt, an evaluation prompt, and an external search algorithm, allows a systematic tree-shaped exploration of the space of reasoning steps. Figure 6 illustrates the different reasoning structures. (Another approach, Graph of Thoughts, allows even more complex relations between the reasoning steps (Besta et al., 2024).)
@@
Many works introduce variants on external prompt-improvement loops, to have explicit control over the reasoning process. They use techniques from planning and tree search to be able to use backtracking to traverse the space of possible combinations of reasoning steps (Yao et al., 2024; Xie et al., 2024; Besta et al., 2024; Schultz et al., 2024; Browne et al., 2012; Gandhi et al., 2024). Other methods are also used for prompt creation. Evolutionary algorithms (Romera-Paredes et al., 2024; van Stein and Bäck, 2024) and planning methods (Bohnet et al., 2024; Valmeekam et al., 2023; Kambhampati et al., 2024) are used to create new prompts and heuristic algorithms for LLMs, and, synergistically, to use LLMs to create new heuristic evolutionary and planning algorithms.
@@
The external search algorithm can also be used to generate training data, for finetuning the LLM, or for pretraining. In this way, we can try to see if an LLM can be taught to search possible steps implicitly, without the need for an external control loop. In the Stream of Search approach Gandhi et al. (2024) create a language for search sequences, and subsequently train an LLM on search trees that contain both good and bad outcomes, improving the accuracy of the model. This approach internalizes the outcome of external searches into the LLM. Schultz et al. (2024) further show how such search results can be used to train an LLM and achieve Grandmaster-level performance in Chess, Connect Four, and Hex.
2.2 Self Reflection

### v3

@@ 1.5. Taxonomy
| Toolformer (schick2023toolformer) | Call APIs | Tool calling | Calculator, Search engine |
| ToolBench (qin2023toolllm) | 16,464 APIs | Tool calling | API framework |
| EasyTool (yuan2024easytool) | Tool documentation | Tool calling | ToolBench |
| ToolAlpaca (tang2023toolalpaca) | 400 APIs | Tool calling | Benchmark |
| ToolQA (zhuang2023toolqa) | APIs | Tool calling | Question answering |
| Gorilla (patil2023gorilla) | Generate APIs | Tool calling | APIBench |
| AgentHarm (andriushchenko2024agentharm) | Adversarial Agents | Robust LLMs | Adversarial Benchmark |
@@
| RainbowTeaming (samvelyan2024rainbow) | MAP Elites | Robust LLMs | Adversarial Benchmark |
@@
| AssistantGPT (neszlenyi2024assistantgpt) | Websearch, OpenAPI, Voice | Tools, Planner, Memory | Education/Corporate |
| Meeting Assist (cabrero2024exploring) | LLM | Meetings | Scrum |
@@
| MUCA (mao2024multi) | topic generator | What/When/How | Group Conversations |
| Task Scheduling (bastola2023llm) | LLM | Task Scheduling | Collaborative Group |
| Thinking Assistant (park2023thinking) | LLM | Human reflection | Human |
| LLaSa (zhang2024large) | finetuned LLM, CoT, RAG | E-commerce assistant | ShopBench |
| MMLU (jin2024shopping) | shopping skills | Finetuning | Benchmark |
| Question suggestion (vedula2024question) | LLM | Product metadata | Shopping |
| ChatShop (chen2024chatshop) | finetuned LLM | Information-seeking | Shopping |
| Flight Booking Assistant (manasa2024towards) | finetuned LLM, RAG | Flight Booking | Booking process |
@@
| Medical Note generation (yuan2024continued) | finetuned LLM | Medical Scribe | Medical note taking |
@@
| Medical Reports (sudarshan2024agentic) | Reflexion (shinn2024reflexion) | 21st Century Cures Act | Health records |
| MedCo (wei2024medco) | Multiagent Copilot | Medical education | education |
| Benchmark (qiao2024benchmarking) | RAG | Agentic Workflow | Benchmark |
| Wind Hazards (tabrizian2024using) | LLM | Flight Planning | Flight Operations |
| Flight Dispatch (wassim2024llm) | LLM | Drone as a Service | Flight Operations |
| FinAgent (zhang2024finagent) | Multimodal, RAG | Analysis modules | Stock data |
| FinRobot (yang2024finrobot) | finetuned LLM | Document Analysis | Financial Documents |
| FinMem (yu2024finmem) | Multi-agent | Trading Agent Assistant | Market data |
| TradingAgents (xiao2024tradingagents) | Multi-agent | Collaborative dynamics | Simulation |
| AI Scientist (lu2024ai) | Chain of Thought | Reflexion (shinn2024reflexion) | Scientific experiment |
| SWE-Agent (yang2024swe) | Codex | ReAct (yao2022react) | Agent-Computer Interface |
| MLGym (nathani2025mlgym) | Chain of Thought | SWE-Agent | Gym (brockman2016openai) |
@@
| Iterated Prisoner’s (fontana2024nicer) | LLM | Cooperate/Defect | Social Dilemma |
@@
| Social Games (akata2023playing) | LLM | Cooperate/Defect | Battle of the Sexes, etc |
@@
| GTBench (duan2024gtbench) | CoT/ToT | Cooperate/Defect | Kuhn poker, liar’s dice, nim |
@@
| GAMA-Bench (huang2024far) | LLM | Cooperate/Defect | El Farol, Public Goods, etc |
| Theory of Mind (van2023theory) | LLM | Theory of Mind | Stories |
| NegotiationArena (bianchi2024well) | LLM | Dialogue | Negotiation |
| Alympics (mao2023alympics) | LLM | Multi-agent sandbox | Water-allocation challenge |
| MAgIC (xu2024magic) | LLM | social interaction | Social Deduction games |
| AucArena (chen2023put) | LLM | Bidding/Goal | Auction |
| EgoSocialArena (hou2024entering) | LLM | Social Intelligence | Cognitive, Situational, Behavioral |
| Donor Game (vallinder2024cultural) | LLM | Reciprocity | Social skill Game |
| Social Simulacra (park2022social) | LLM | Society | Simulation of Society, Party |
| Reconcile (chen2023reconcile) | LLM | Concensus | Round Table Conference |
@@
| MindStorms (zhuge2023mindstorms) | LLM | Society of Mind (minsky1988society) | Multi-agent problem solving |
| AutoGen (wu2023autogen) | LLM infrastructure | aegnt-agent conversation | Framework |
| AgentVerse (chen2023agentverse) | LLM | Group dynamics | Collaborative problem solving |
| ChatEval (chan2023chateval) | LLM | Collaborative problem solving | Text summarization |
| CAMEL (li2023camel) | LLM infrastructure | Multi-agent interaction | Roleplaying Framework |
| OASIS (yang2024oasis) | lightweight LLM | Social media simulator | Reddit/X |
| WebArena (zhou2023webarena) | Web benchmark | e-commerce, forum, content | Benchmark |
| Balrog (paglieri2024balrog) | RL games | interaction | Benchmark |
| BenchAgents (butt2024benchagents) | Planning | human in the loop | Benchmark |
| AgentBoard (ma2024agentboard) | Embodied, Web, Tool | interactions | Benchmark |
| Bias (fernando2024quantifying) | LLM | healthcare, justice, business | Benchmark |
@@
| Citing (feng2023citing) | Curriculum Learning | Teacher/Student | Instruction Tuning |
| WEBRL (qi2024webrl) | Curriculum Learning | Self-evolving | WebArena |
@@
| Expert Iteration (zhao2024automatic) | Curriculum Learning | Reasoning | Hallucination Mitigation |
| EvolutionaryAgent (li2024agent) | Evolutionary LLM | Norm Aligment | Multi-agent Infrastructure |
@@
| Social Conventions (ashery2024dynamics) | Naming Game | Norm emergence | Naming game (steels1995self) |
| MetaNorms (horiguchi2024evolution) | LLM | Norm emergence | Metanorms (axelrod1986evolutionary) |
| Norm Violations (he2024norm) | LLM | Norm violations | 80 household stories |
| CASA (qiu2024evaluating) | LLM | Cultural and Social awareness | Benchmark |
| Collaboration (zhang2023exploring) | LLM | 4-traits, cooperation | LLM societies |
@@
| Power hierarchy (campedelli2024want) | LLM | persuasive/abusive behavior | Stanf Prison Exper (zimbardo1972stanford) |
| Argumentation (van2024hybrid) | Hybrid LLM | LLM supported Argumentation | Benchmark |
@@
| Debate (baltaji2024conformity) | LLM-agents | collaboration, debate | Multi-agent discussion |
2. Reasoning
@@
Note that both retrieval augmentation and self reflection can be used to generate new training data. Retrieval augmentation can be used to retrieve relevant information beyond the originally available training dataset. Self reflection uses methods related to planning that imagine plausible futures, that can be useful for training of LLMs. Originally, the methods that we review in this section were developed with the goal of improving the predictive modeling performance of the LLM. For the field of agentic LLMs, the reasoning techniques are used as an important fundament for agents that act with the world, and interact with each other.
We will start with a survey of the individual approaches. Reasoning methods are the foundation of Agentic LLM. In Section 2.4 we will discuss two essential approaches, Chain of Thought and Self Reflection, in more detail.
2.1. Multi Step Reasoning
@@
We will start by reviewing works that apply reasoning methods to improve decision making, inspired by Chain of Thought’s step-by-step approach (wei2022chain).
2.1.1. Chain of Thought Step-by-Step
@@
Originally, LLMs performed poorly on math word problems, even on simple grade school problems (GSM8K, cobbe2021training). LLMs are trained to produce an immediate answer to a prompt, and they typically take shortcuts that may look good, but are semantically wrong.222What is the correct answer to: This is as simple as two minus two is …? The phrase: as simple as two plus two is four may well have a higher frequency in a training corpus than the phrase: as simple as two minus two is zero.
@@
To correctly solve complex reasoning problems, humans are taught to use a step-by-step approach. If a reasoning problem is better solved by following a step-by-step approach, then a sensible approach is to prompt the model to follow suitable intermediate steps, answer those, and work towards the final answer. wei2022chain showed in their Chain of Thought paper that with the right prompt the LLM follows such intermediate steps. When the LLM is prompted to first rephrase information from the question as intermediate reasoning steps in its answer, the LLM performed much better than when it was prompted to answer a math problem directly, without reproducing the information from the question in its answer (see their example in Figure 4).
@@
kojima2022large find that the addition of a single standard phrase to the prompt (Let’s think step by step) already significantly improves performance. Chain of Thought prompts have been shown to significantly improve performance on benchmarks that included arithmetic, symbolic, and logical reasoning.
@@
Long reasoning chains, however, introduce a challenge, since with more steps hallucination increases. A verification method is needed to prevent error-accumulation. A popular approach is Self Consistency (wang2022self). Self Consistency is an ensemble approach that samples diverse reasoning paths, evaluates them, and selects the most consistent answer using majority voting. It improves the performance of Chain of Thought typically by 10-20 percentage points when tested on benchmarks. Prompt improvement approaches based on Chain of Thought and Self Consistency are being used to train most modern reasoning LLMs, including OpenAI o1, o3, DeepSeek and Qwen (wu2024comparative; guo2025deepseek; yang2025qwen3).
2.1.2. Interpreter and Debugger
@@
To solve problems that require mathematical or formal reasoning, it is often advantageous to reformulate the problem into a mathematical or programming language. This reformulated problem can then be solved by a specialized system, such as a mathematical reasoner (moura2021lean), an interpreter, or a planner.
@@
LLMs are not just successful in natural languages, but also in formal (computer) languages. Codex is an LLM that is pretrained on computer programs from GitHub (chen2021evaluating), which has been successfully deployed commercially. Codex has been used as the basis for the MathPrompter system (imani2023mathprompter). MathPrompter is an ensemble approach that generates algebraic expressions or Python codes, that are then solved using a math solver, or a Python interpreter. Using this approach, MathPrompter achieves state-of-the-art results on the MultiArith dataset (from 78.7% to 92.5%), with GPT-3.
@@
Two other approaches that use a formal language are Program of Thought (PoT) (chen2022program) and Program Aided Language (PAL) (gao2023pal). Both approaches generate Python code and use the Python interpreter to evaluate the result.
@@
Debuggers can be used to provide feedback on generated code. This approach is followed in the Self Debugging work (chen2023teaching), that teaches an LLM to debug its generated program code. It follows the same steps of code generation, code execution, and code explanation that a human programmer follows. Several works use Self Debugging to generate code tuned for solving specific problems automatically, without human feedback. romera2024mathematical introduced FunSearch, an approach that integrates formal methods and LLMs to enhance mathematical reasoning and code generation. It uses a genetic algorithm approach with multiple populations of candidate solutions (programs), which are automatically evaluated (using tools depending on the problem specification). LLaMEA (Large Language Model Evolutionary Algorithm) leverages evolutionary computation methods to generate and optimize evolutionary algorithms (van2024llamea).
@@
Planners are also combined with LLMs at the language level. bohnet2024exploring provide a benchmark for PDDL (aeronautiques1998pddl) based planning problems. They study how LLMs can achieve success in the planning domain (Figure 5).
@@
In Section 2.3 on retrieval augmentation, we will see further approaches where deep learning and symbolic approaches are successfully combined (gao2023pal).
2.1.3. Search Tree
@@
Chain of Thought uses a prompt that causes the model to perform a sequence of steps. When there is a single next step, that will be taken. When there are more possibilities, it is unclear how the next step should be selected. A greedy method selects the single step that looks best, follows only that step, and forgets the alternatives (Chain of Thought). Ideally, we should follow the tree of all possible steps. This method is chosen in the Tree of Thoughts approach (yao2024tree). Here, an external control algorithm is created, that calls the model, each time with a different prompt, so that it follows a tree of reasoning steps. When one reasoning path has been traversed, the search backtracks, and tries an alternative. The paper describes both a breadth-first and a depth-first controller.
@@
Together, the trio that consists of a generation prompt, an evaluation prompt, and an external search algorithm, allows a systematic tree-shaped exploration of the space of reasoning steps. Figure 6 illustrates the different reasoning structures. (Another approach, Graph of Thoughts, allows even more complex relations between the reasoning steps (besta2024graph).)
@@
Many works introduce variants on external prompt-improvement loops, to have explicit control over the reasoning process. They use techniques from planning and tree search (hart1968formal; plaat1996research) to be able to use backtracking to traverse the space of possible combinations of reasoning steps (yao2024tree; xie2024self; besta2024graph; schultz2024mastering; browne2012survey; gandhi2024stream). Other methods are also used for prompt creation. Evolutionary algorithms (romera2024mathematical; van2024llamea) and planning methods (bohnet2024exploring; valmeekam2023planning; kambhampati2024llms) are used to create new prompts and heuristic algorithms for LLMs, and, synergistically, to use LLMs to create new heuristic evolutionary and planning algorithms.
@@
The external search algorithm can also be used to generate training data, for finetuning the LLM, or for pretraining. In this way, we can try to see if an LLM can be taught to search possible steps implicitly, without the need for an external control loop. In the Stream of Search approach gandhi2024stream create a language for search sequences, and subsequently train an LLM on search trees that contain both good and bad outcomes, improving the accuracy of the model. This approach internalizes the outcome of external searches into the LLM. schultz2024mastering further show how such search results can be used to train an LLM and achieve Grandmaster-level performance in Chess, Connect Four, and Hex.
2.2. Self Reflection
Reasoning methods draw inspiration from step-by-step human solution approaches. The more elaborate approaches use explicit planning-like methods to look ahead and use feedback for verification. These methods use a form of reinforcement learning, the type of machine learning where the agent learns a policy of actions to take from reward feedback from the environment states (see Figure 7 (sutton2018reinforcement; plaat2022deep)). Such prompt-improvement loops facilitate a form of self reflection, since the model assesses and improves its own. In reinforcement learning terms, both the agent and the environment are the LLM, but with different prompts.